OpenStudy (anonymous):

Write the slope-intercept equation for the line through (3, 7) and (9, 3).

OpenStudy (hoblos):

slope = variation of y over variation of x = (3-7)/(9-3) = -4/6 = -2/3 y-7=-2/3(x-3) y= (-2/3)x +9

OpenStudy (anonymous):

your awesome

OpenStudy (hoblos):

:D